A galvanic cell is set up from a zinc bar weighing 50 gram and 1.0 litre of 1.0 molar CuSO4 solution. How long would the cell run, assuming  it delivers a steady current of 1.0 ampere
(a) 48 hrs (b) 41 hrs (ć) 21 hrs (d) 1hrs

Grade:12th pass

1 Answers

Arun
5 years ago

As u know According to Faraday's 1st law if electrolysis….

W=Z×q or W=Z×I×t

We know Z ( electrochemical equivalent):—

Z= Equivalent weight / Faraday const. (F)

So, Equivalent wt of CuSO4 = molecular mass of CuSO4 valency.

Z= 159/2 = 79.5

Now we r given with,

W= 50g

I=1A

t=???

So put values in Formula and time will be founded…